• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name solvignon

    Log entry time 17:37:01 on June 2,2002

    Entry number 85116

    keyword=target rebooted

    We lost the signal from hydrogen and deuterium charts but not from coolant.

    So rebooted IOC and turned on the manual heater. But target temperature
    kept falling down and the signal didn't come back. So rebooted again.
    After 2 minutes the PID worked again and the targets came back to the right
    temperature.

    It looks that the manual heater is not working properly so one will have to
    test it when beam will be down and make a Halog entry when done.

    The test consists of turning off the PID and turning on the manual heater.
    If the target temperature keeps falling down, then the manual heater is not
    working and we will have to tell the target expert. But don't let the temperature
    going below 19 K.